OpenStudy (anonymous):

Find the sum of all real solutions of the equation x^2−16=3(x−4)^2.

OpenStudy (anonymous):

Solve for x over the real numbers: x^2-16 = 3 (x-4)^2 Write the quadratic polynomial on the right hand side in standard form. Expand out terms of the right hand side: x^2-16 = 3 x^2-24 x+48 Move everything to the left hand side. Subtract 3 x^2-24 x+48 from both sides: -2 x^2+24 x-64 = 0 Factor the left hand side. The left hand side factors into a product with three terms: -2 (x-8) (x-4) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by -2: (x-8) (x-4) = 0 Solve each term in the product separately. Split into two equations: x-8 = 0 or x-4 = 0 Look at the first equation: Solve for x. Add 8 to both sides: x = 8 or x-4 = 0 Look at the second equation: Solve for x. Add 4 to both sides: Answer: | | x = 8 or x = 4

OpenStudy (anonymous):

\[\left( a-b \right)^{2}=a ^{2}+b ^{2}-2ab\]

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\] sum of roots=-b/a

OpenStudy (anonymous):

so the answer is 4 ?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

x^2-16=3(x-4)^2 x^2-16=3(x^2-8x+16) 2x^2-24x+64=0 x^2-12x+32=0 sum of roots=-(-12)/1=12